# Feature-Based Interpretable Optimization

### v1

@@ 3 Models
In this section, we discuss how we can obtain decision tree-based optimization rules using a mixed-integer programming (MIP) formulation and how it can be used to obtain the meta-solutions for knapsack and shortest path problems.
@@ 3.1 Inner Tree Structure
We start by introducing an MIP formulation for the generation of decision trees which is part of the models in the following sections. The formulation presented is taken from [GH23]. It can be used to generate univariate binary decision trees of fixed depth $Q$. The presented constraints only govern the inner structure of the tree, i.e., the branching decisions and the assignment of data points to the leaves. We also need additional constraints to assign solutions spaces (or specific solutions) as well as an objective function that captures the selected decision criteria. Various extensions and variations of this model can be found in [GH23], which are compatible with the formulations presented in this paper.
For simplicity, we present the formulation for symmetric decision trees. Such trees require all the queries at any given depth $q$ of the tree to be the same. For every given data point $j$ and every leaf $k$, a binary variable $\ell_{j}^{k}$ indicates if the data point $j$ reaches leaf $k$ following the tree’s queries. Each query $q$ is modeled using a binary variable $d_{f}^{q}$, indicating if feature $f$ is evaluated in $q$ and a continuous variable $b^{q}$ specifying its threshold. The latter can be bounded by the minimum and the maximum of all feature values, see (1f). Via the constraints (1b) it is ensured that every data point from our training set is assigned to exactly one leaf of the tree. The constraints (1c) enforce that for every query in the tree, a feature has to be chosen. Constraints (1d) and (1e) control the assignment of data points to leaves while taking into account the relevant tree queries. A data point can reach the left child if and only if the value of instance feature $i$ of data point $j$ is smaller than or equal to the chosen threshold for this query, see (1d). Similarly, if the value is greater than the threshold it can reach the right child, see (1e). This makes use of index sets $S_{q}$ to control the assignment of left/right paths to leaves (for details, we refer to [GH23]).
Note that any other formulation for decision trees can be used for the following models if it includes variables $\ell^{k}_{j}$ which indicate the mapping of a scenario $j$ to a leaf $k$.
[…]
@@
$\sum_{f\in[F]}d_{f}^{q}=1$ $\forall q\in[Q]$ (1c)
$\sum_{f\in[F]}\phi_{I}^{jf}d_{f}^{q}\leq b^{q}+M\sum_{k\in S_{q}}\ell^{k}_{j}$ $\forall q\in[Q],\ j\in[N]$ (1d)
$b^{q}+\epsilon-M(1-\sum_{k\in S_{q}}\ell^{k}_{j})\leq\sum_{f\in[F]}\phi_{I}^{jf}d_{f}^{q}$ $\forall q\in[Q],\ j\in[N]$ (1e)
$b^{q}\in[\min_{j\in[N],f\in[F]}\phi_{I}^{jf}-\epsilon,\max_{j\in[N],f\in[F]}\phi_{I}^{jf}]$ $\forall q\in[Q]$ (1f)
[…]
3.2 Interpretable Knapsack Problems
We return to the knapsack problem from Section 2.2. In this setting, the set of items can be fragmented into several subsets. For example, when packing a bag for a vacation there could be a set of shirts, sweaters, and jackets. Let us say the optimization rule $a$ takes in the parameters of the instance (things like weights, values, and capacity) and via a small decision tree yields characteristics for each subproblem. E.g., if it is a one-week summer vacation on the beach pack 7 shirts, 1 sweater, and 1 jacket, while you should pack 4 shirts, 3 sweaters, and 2 jackets for an autumn hiking trip. We assume that the knapsack capacity $C\in\mathbb{R}$ as well as the item weights $w_{i}\in\mathbb{R}$ are fixed. Then, finding a suitable rule $a$ implies finding an appropriate decision tree that can specify the capacities for the different item groups given the problem instances. Using the Laplace decision criterion and assuming a set of $N$ historic observations this can be achieved via solving the following optimization problem.
$\max$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{j\in[N]}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{c}_{j}^{\top}\bm{x}_{j}$ (2a)
s.t. $\bm{\ell}\in T(\phi_{I}^{1},\ldots,\phi_{I}^{N})$ (2b)
$\sum_{i\in I_{f}}w_{i}x_{ji}\leq C^{k}_{f}+M(1-\ell_{j}^{k})$ $\forall f\in[F],\ k\in[K],\ j\in[N]$ (2c)
@@ 3.2 Interpretable Knapsack Problems
3.3 Interpretable Shortest Path Problems
3.3.1 MIP Formulation
In this section, we describe the formalization of the shortest path approach sketched in Section 2.2. Like the knapsack formulation, we also use the Laplace criterion in the shortest path objective. In contrast to the setting for the knapsack model, however, formalizing this problem is not so straightforward. The difficulty lies in describing the solution space in the formulation. We want to ensure that every feasible $s$-$t$-path has to follow its given meta-solution. That means that the path must visit the districts (i.e., at least one node of every given district) specified in the meta-solution (also referred to as the meta-path) in the specified order. The corresponding district of a node $v$ is indicated in the following with ${f}(v)$.
@@ 3.3.1 MIP Formulation
Given a graph $G=(E,V)$ an auxiliary graph $G^{\prime}=(V^{\prime},E^{\prime})$ is used to formulate this approach as an IP. The set of nodes is given by
@@
where $V_{\delta}$ represents the $\delta$th copy of $V$. For $\delta\in[\Delta]$ let $v_{\delta}$ be the node of $V_{\delta}$ that is a copy of $v\in V$. The sets $V_{\delta}$ are, in the following, referred to as layers. The set of edges is given by
$E^{\prime}=\bigcup_{(u,w)\in E:{f}(u)={f}(w)}\bigcup_{\delta\in[\Delta]}\{(u_{\delta},w_{\delta})\}\cup\bigcup_{(u,w)\in E:{f}(u)\neq{f}(w)}\bigcup_{\delta\in[\Delta-1]}\{(u_{\delta},w_{\delta+1})\},$
@@
We now consider the problem of finding the shortest path from $s_{0}$ to some $t_{\delta}$ in $G^{\prime}$ with the restriction that for every $v\in V$ we are only allowed to visit at most one of its copies $v_{\delta}\in V^{\prime}$. Further, we only allow an edge $(u_{\delta},w_{\delta+1})\in E^{\prime}$ connecting different layers to be used if the corresponding node $u\in V$ is located in the district which is assigned to layer $\delta$ and the corresponding node $w\in V$ is located in the district which is assigned to layer $\delta+1$. Given a meta-solution and a $s_{0}$-$t_{\delta}$-path for some $\delta\in\Delta$, both represented in $G^{\prime}$, we can obtain a $s$-$t$-path in $G$ which follows the meta-solution.
[…]
@@
Figure 6: $G$ and corresponding $G^{\prime}$ with an exemplary meta-solution highlighted.
We are using the sets for formalizing this idea in an MIP formulation:
$E^{v\delta}_{-\rightarrow}$ =$=\{(u_{\delta-1},v_{\delta})\in E^{\prime}\},$
$E^{v\delta}_{-\downarrow}$ =$=\{(u_{\delta},v_{\delta})\in E^{\prime}\},$
$E^{v\delta}_{+}$ =$=\{(v_{\delta},w_{\delta})\in E^{\prime}\}\cup\{(v_{\delta},w_{\delta+1})\in E^{\prime}\},$
$E_{\rightarrow}$ $=\cup_{v\in V}\cup_{\delta\in\Delta}E_{-\rightarrow}^{v\delta}.$
For the following, we consider not only edges but also nodes as belonging to layers. It applies that edges within a layer $(u_{\delta},w_{\delta})$ and edges which leave it ($u_{\delta},w_{\delta+1}$) are considered belonging to this layer $\delta$. For the calculation of the objective value and for ensuring that every meta-solution offers the possibility of finding an $s$-$t$-path, for every scenario $j$ a solution $\bm{x}^{\delta}_{j}$ has to be calculated. The binary variables $x_{ej}^{\delta}$ indicate when set to one that an edge $e\in\{(u_{\delta},w_{\delta})\cup(u_{\delta},w_{\delta+1})\in E^{\prime}:\delta\in[\Delta]\}$ being in or leaving layer $\delta$ is part of the solution of scenario $j$. Constraints (3c) to (3e) represent flow conservation constraints, ensuring that $\bm{x}^{\delta}_{j}$ represents a $s_{0}$-$t_{\delta}$-path for some $\delta\in[\Delta]$ in $G^{\prime}$. Since every node $v\in V$ has multiple representants in $V^{\prime}$ it is necessary to ensure that a path found in $G^{\prime}$ does not contain cycles in $G$. Constraints (3f) allows that at most one representant of any node $v\in V$ can be part of a path in $G^{\prime}$. The assignment of these variables will be more restricted in the following.
As seen, it is essential to couple the district/layer assignment to the edges traversing between the layers. To control these edges, we are using binary variables $y^{\delta fk}$. Edges within one layer are not bounded by them. When set to one, $y^{\delta fk}$ indicates that feature/district $f$ is assigned to layer $\delta$ in leaf $k$. Constraints (3g) ensure that at most one feature is assigned to every layer $\delta$ in every leaf $k$. The constraints (3h) restrict the usage of edges between layers for every scenario that is assigned to leaf $k$. An edge $(u_{\delta},w_{\delta+1})$ connecting two consecutive layers can only be used if $y^{\delta fk}=1:f={f}(u)$ and $y^{(\delta+1)fk}=1:f={f}(w)$. Those edges can only be used if the districts assigned to the start and end layers match the features of the start and end nodes of that edge. Constraints (3i) and (3j) ensure that the sink nodes $t_{\delta}$ can only visited in layer $\Delta$ or if its last layer has no feature assigned, indicating the end of solution sequence.
Note that if there are leaves that have no assigned scenarios, the formulation does not enforce it to contain a feasible solution. In practice, if such a case happens, the tree can either be pruned in post-processing, or a random feasible meta-solution can be placed in this leaf.
$\min\quad$ $\sum_{j\in[N]}\sum_{\delta\in\Delta}\bm{c}_{j}^{\top}\bm{x}_{j}^{\delta}$ (3a)
s.t. $\bm{\ell}\in T(\phi_{I}^{1},\ldots,\phi_{I}^{N})$ (3b)
@@
$\sum_{f\in[F]}y^{\delta fk}\leq 1$ $\forall\delta\in[\Delta],k\in[K]$ (3g)
@@
$x_{j,e}^{\delta-1}\leq(1-\sum_{f\in[F]}y^{(\delta+1)fk})+(1-\ell_{j}^{k})$ $\forall\delta\in[\Delta-1],k\in[K],j\in[N],e\in E_{-\rightarrow}^{t\delta}$ (3i)
$x_{j,e}^{\delta}\leq(1-\sum_{f\in[F]}y^{(\delta+1)fk})+(1-\ell_{j}^{k})$ $\forall\delta\in[\Delta-1],k\in[K],j\in[N],e\in E_{-\downarrow}^{t\delta}$ (3j)
@@
$y^{\delta fk}\in\{0,1\}$ $\forall\delta\in[\Delta],f\in[F],k\in[K].$ (3l)
3.3.2 Hardness Results
We now turn to the complexity of the interpretable shortest path problem. As it is possible to solve the nominal shortest path problem in graphs without negative cycles efficiently, the question arises whether polynomial-time algorithms also exist for the interpretable case. In the following, we give negative answers to two special cases.
Theorem 1.
Proof.
A reduction from the NP-complete Hamiltonian path problem [GJ79] is constructed. Given an undirected graph $G=(V,E)$ with nodes $s,t\in V$, we need to decide whether there is a (simple) $s$-$t$-path visiting all nodes in $V$ exactly once. We first create a graph $\bar{G}=(\bar{V},\bar{E})$ in the following way: For every $v\in V$, two nodes $v$ and $v^{\prime}$ are added to $\bar{G}$ as well as an edge $(v,v^{\prime})$. Node $v$ is assigned to district $a$, and node $v^{\prime}$ to district $b$. For every $\{i,j\}\in E$, the edges $(i^{\prime},j)$ and $(j^{\prime},i)$ are added to $\bar{G}$. Figure 7 shows an example construction.
Figure 7: Example construction for the proof of Theorem 1.
Theorem 2.
The following problem is NP-complete: Given a graph $G=(V,E)$ where each node $v\in V$ belongs to a district $f(v)$, and a list of cost scenarios $\bm{c}^{1},\ldots,\bm{c}^{N}\in\{0,1\}^{E}$. Decide whether there is a meta-path such that the corresponding paths in each scenario have costs 0. In particular, the problem of finding a single meta-path with minimum costs is not approximable, even in directed acyclic graphs.
Proof.
Let an instance of the NP-complete 3SAT problem [GJ79] be given. It consists of $n$ Boolean variables $x_{1},\ldots,x_{n}$ and a list of clauses $C_{1},\ldots,C_{m}$, where each clause is a disjunction of three literals. We need to decide if there is a truth assignment of the $n$ variables such that all $m$ clauses are true.
We first show how to construct the graph $G^{\prime}$ that shows only one node for all nodes in $G$ that belong to the same region. It consists of $22n+$ 2 nodes $V^{\prime}$. For each variable $x_{i}$, we construct a node $t_{i}$ and a node $f_{i}$, which represent the states of “true” and “false”, respectively. We further introduce nodes $s$ and $t$ to be used as the source and target nodes of the meta-path. The set of edges is given by
$\cup\left\{(s,t_{1}),(s,f_{1}),(t_{n},t),(f_{n},t)\right\}.$
$E=$ $E_{1}\cup E_{2}$
$\cup\{(t^{1}_{i},t^{2}_{i}):i\in[n]\}\cup\{(f^{1}_{i},f^{2}_{i}):i\in[n]\}$
$\cup\{(s,t^{1}_{1}),(s,f^{1}_{1}),(t^{2}_{n},t),f^{2}_{n},t)\}$
with $\text{with }E_{1}=$$\Big{\{}(a,b):a\in\{t^{1}_{i},f^{1}_{i}\},\ b\in\{t^{1}_{i+1},f^{1}_{i+1}\},\ i=1,\ldots,n-1\Big{\}}$
Finally, we introduce $m$ cost vectors $\bm{c}^{j}\in\{0,1\}^{E}$, each of them corresponding to a clause $C_{j}$. The costs for all edges in $E_{1}$ and $E_{2}$ are always zero. The costs of an edge $(t^{1}_{i},t^{2}_{i})$ is zero if $x_{i}$ is contained in $C_{j}$, otherwise, the costs are one. Analogously, the costs of an edge $(f^{1}_{i},f^{2}_{i})$ is zero if $\bar{x}_{i}$ is contained in $C_{k}$, otherwise, the costs are one.
Finally, the costs of edges $(s,t^{1}_{1})$, $(s,f^{1}_{1})$, $(t^{2}_{n},t)$, and $(f^{2}_{n},t)$ are always zero.
Figure 9 illustrates the construction of $G$ with the cost scenario that corresponds to the clause $(x_{1}\vee\bar{x}_{2}\vee x_{3})$. There are only three edges with costs one, indicated as a dotted arrow. These are the edges $(f^{1}_{1},f^{2}_{1})$, $(t^{1}_{2},t^{2}_{2})$, and $(f^{1}_{3},f^{2}_{3})$.
Figure 9: Graph in proof of Theorem 2.

### v3

@@ 3 Models
In this section, we discuss how we can obtain decision tree-based optimization rules using a mixed-integer programming (MIP) formulation and how it can be used to obtain the meta-solutions for knapsack and shortest path problems. Appendix B includes hardness proofs for two special cases of shortest path problems and also highlights the appropriateness of MIP as a modelling tool.
@@ 3.1 Inner Tree Structure
We start by introducing an MIP formulation for the generation of decision trees which is part of the models in the following sections. The formulation presented is adapted from Goerigk and Hartisch (2023). It can be used to generate univariate binary decision trees of fixed depth $Q$. The presented constraints only govern the inner structure of the tree, i.e., the branching decisions and the assignment of data points to the leaves. We also need additional feasibility constraints for the solutions assigned to the tree leaves as well as an objective function that captures the selected decision criteria. Various extensions of this model can be found in the literature, which are compatible with the formulations presented in this paper.
For simplicity, we present the formulation for symmetric decision trees of depth $Q$. Such trees require that all queries at any given depth $q\in[Q]$ of the tree be the same. For every given data point $j$ and every leaf $k$, a binary variable $\ell_{j}^{k}$ indicates if the data point $j$ reaches leaf $k$ following the tree’s queries. Each query $q\in[Q]$ is modeled using a binary variable $d_{f}^{q}$ which indicates whether feature $f$ is queried at depth $q$. A continuous variable $b^{q}$ specifies the threshold for the evaluation. It can be bounded by the minimum and the maximum of all feature values, see (1f). The Constraints (1b) ensure that every data point from the training set is assigned to exactly one leaf of the tree. To ensure that every query in the tree chooses a feature, we enforce the Constraints (1c). Constraints (1d) and (1e) control the assignment of data points to leaves while taking into account the relevant tree queries. A data point $j$ can reach the left child of query $q$ if and only if the value of the instance feature $f$ of data point $j$ is smaller than or equal to the chosen threshold for this query, see (1d). Similarly, if the value is greater than the threshold, it can reach the right child, see (1e). To this end, we make use of the index sets $S_{q}$ to control the assignment of left/right paths to leaves. For a given query node $q$, $S_{q}\subseteq[K]$ contains the leaves which can be reached by traversing to the right child of $q$. E.g. for $Q:=2$ and therefore $K=4$ we have $S_{1}=\{3,4\},S_{2}=\{2,4\}$.
Note that other formulations for decision trees can also be used in the upcoming models as long as they include the variables $\ell^{k}_{j}$, which indicate a mapping of a scenario $j$ to a leaf $k$.
$T(\bm{\phi}_{I}^{1},\ldots,\bm{\phi}_{I}^{N})=\Big\{$ $\bm{\ell}\in\{0,1\}^{K\times N}\text{ s.t. }\exists\bm{d},\bm{b}:$ (1a)
@@
$\sum_{f\in[F_{I}]}d_{f}^{q}=1$ $\forall q\in[Q]$ (1c)
$\sum_{f\in[F_{I}]}\phi^{jf}_{I}d_{f}^{q}\leq b^{q}+M\sum_{k\in S_{q}}\ell^{k}_{j}$ $\forall q\in[Q],\ j\in[N]$ (1d)
$b^{q}+\epsilon-M(1-\sum_{k\in S_{q}}\ell^{k}_{j})\leq\sum_{f\in[F_{I}]}\phi^{jf}_{I}d_{f}^{q}$ $\forall q\in[Q],\ j\in[N]$ (1e)
$b^{q}\in[\min_{j\in[N],f\in[F_{I}]}\phi^{jf}_{I}-\epsilon,\max_{j\in[N],f\in[F_{I}]}\phi^{jf}_{I}]$ $\forall q\in[Q]$ (1f)
$d_{f}^{q}\in\{0,1\}$ $\forall f\in[F_{I}],\ q\in[Q]$ $\Big\}.$ (1g)
3.2 Interpretable Surrogates for the Knapsack Problem
We return to the knapsack problem from Section 2.2. In this setting, the set of items can be categorized into several subsets. For example, when deciding about in-house budget allocation, the relevant projects can be grouped by their purpose—likewise to Chao and Kavadias (2008). There could be a set of projects affiliated with marketing, production, or research and development. We assume their investment costs to be known, but their pay-off are in the future and hence uncertain. Let us say the optimization rule $a$ takes in the features of the instance (e.g., cost, expected pay-off, demand, macroeconomic factors) and via a small decision tree yields the meta-solutions, i.e., it specifies the budgets for the different categories. For example, in times of high demand, the proportion of budget allocated to production-affiliated projects can be increased. In contrast, whilst facing a low demand, it is desirable to shift a higher share of the total budget towards marketing efforts. Whilst these decision rules are created in agreement with the strategic management of the company, choosing the specific projects to execute is the responsibility of an operational-level manager facing the observed realization.
We assume that the knapsack capacity $C\in\mathbb{R}$ as well as the item weights $w_{i}\in\mathbb{R}$ are fixed. Using the Laplace decision criterion and assuming a set of $N$ historic observations, this can be achieved via solving the following optimization problem.
$\max\$$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{j\in[N]}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{c}_{j}^{\top}\bm{x}_{j}$ (2a)
s.t. $\bm{\ell}\in T(\bm{\phi}_{I}^{1},\ldots,\bm{\phi}_{I}^{N})$ (2b)
$\sum_{i\in I_{f}}w_{i}x_{ji}\leq C^{k}_{f}+M(1-\ell_{j}^{k})$ $\forall f\in[F_{I}],\ k\in[K],\ j\in[N]$ (2c)
@@ 3.2 Interpretable Surrogates for the Knapsack Problem
3.3 Interpretable Surrogates for the Shortest Path Problem
In this section, we describe the formalization of the shortest path approach sketched in Section 2.2. For the shortest path problem, each scenario or instance feature corresponds to a different cost vector. The rest of the problem, such as the feasible region and the source and target, remains the same across all the instances. As described in Section 2.2, each node is associated with a district, and the solution features represent a path between the source and target, consisting of a series of districts. Similar to the knapsack formulation, we also use the Laplace criterion in the shortest path objective.
In contrast to the knapsack model, formalizing the shortest path problem is not so straightforward. The difficulty lies in describing the solution space in the formulation. We want to ensure that every feasible $s$-$t$-path has to follow its given meta-solution. That means that the path must visit the districts (i.e., at least one node of the district) specified in the meta-solution (also referred to as the meta-path) in the specified order. We denote the district containing any node $v$ by f(v){f}(v).
@@ 3.3 Interpretable Surrogates for the Shortest Path Problem
Given a graph $G=(E,V)$, we can construct an auxiliary graph $G^{\prime}=(V^{\prime},E^{\prime})$ to formulate the meta-solution problem as an IP. The set of nodes is given by
@@
[…]
$E^{\prime}=\left(\bigcup_{(u,w)\in E:{f}(u)={f}(w)}\;\bigcup_{\delta\in[\Delta]}\{(u_{\delta},w_{\delta})\}\right)\cup\left(\bigcup_{(u,w)\in E:{f}(u)\neq{f}(w)}\;\bigcup_{\delta\in[\Delta-1]}\{(u_{\delta},w_{\delta+1})\}\right),$
@@
We now consider the problem of finding the shortest path from $s_{0}$ to some $t_{\delta}$ in $G^{\prime}$ with the restriction that for every $v\in V$ we are only allowed to visit at most one of its copies $v_{\delta}\in V^{\prime}$. The necessity of explicitly enforcing this condition is proven in Appendix A. Further, we only allow an edge $(u_{\delta},w_{\delta+1})\in E^{\prime}$ connecting different layers to be used if the corresponding node $u\in V$ is located in the district which is assigned to layer $\delta$ and the corresponding node $w\in V$ is located in the district which is assigned to layer $\delta+1$. Given a meta-solution and an $s_{0}$-$t_{\delta}$-path for some $\delta\in\Delta$, both represented in $G^{\prime}$, we can obtain an $s$-$t$-path in $G$ which follows the meta-solution.
Figure 7 shows an example graph $G$ and the corresponding graph $G^{\prime}$ with $\Delta:=6$. Colors are used to represent the district containing the nodes. Assume that we are given a meta-solution represented by the sequence (red-green-yellow-green-pink). Possible resulting $s$-$t$-paths in $G$ are (s(s-11-33-44-55-$t)$ and (s(s-33-44-55-$t)$, which correspond to the two paths in $G^{\prime}$. It can be seen that it is ensured that at least one node in every district listed in the meta-solution has to be visited. The construction also ensures that the order specified by the sequence is adhered to. Additionally, it is not necessary to assign a district to every layer, as long as the sequence of assigned districts allows for the reachability of a node $t_{\delta}$ for some $\delta\in[\Delta]$.
@@
Figure 7: $G$ and corresponding $G^{\prime}$ with an exemplary meta-solution indicated. Edges in $G^{\prime}$ highlighted in black represent edges that can be used for finding a path from $s_{1}$ to some $t_{\delta}$ within the solution space defined by this meta-solution. Grey edges may not be used.
We use the following sets to formalize this idea in an MIP formulation:
  • $E^{v\delta}_{-\rightarrow}=\{(u_{\delta-1},v_{\delta})\in E^{\prime}\}:$ edges coming into node $v$ in layer $\delta$ from layer $\delta-1$,
  • $E^{v\delta}_{-\downarrow}=\{(u_{\delta},v_{\delta})\in E^{\prime}\}:$ edges coming into node $v$ from within layer $\delta$,
  • $E^{v\delta}_{+}=\{(v_{\delta},w_{\delta})\in E^{\prime}\}\cup\{(v_{\delta},w_{\delta+1})\in E^{\prime}\}:$ edges going out from node $v$ within layer $\delta$ and to layer $\delta+1$,
  • $E_{\rightarrow}=\cup_{v\in V}\cup_{\delta\in\Delta}E_{-\rightarrow}^{v\delta}:$ the set of all edges coming into any layer from previous layers.
For modeling, we consider edges as belonging to particular layers. Specifically, edges within a layer $(u_{\delta},w_{\delta})$ and edges which leave it ($u_{\delta},w_{\delta+1}$) are considered to belong to the layer $\delta$. For the calculation of the objective value and to ensure that every meta-solution offers the possibility of finding an $s$-$t$-path, we calculate a solution $\bm{x}^{\delta}_{j}$ for every scenario $j$. The binary variable $x_{ej}^{\delta}$ is equal to $1$ if the edge $e\in\{(u_{\delta},w_{\delta})\cup(u_{\delta},w_{\delta+1})\in E^{\prime}:\delta\in[\Delta]\}$ belonging to layer $\delta$ is part of the solution of scenario $j$. Constraints (3c) to (3e) represent flow conservation constraints, ensuring that $\bm{x}^{\delta}_{j}$ represents a valid $s_{0}$-$t_{\delta}$-path for some $\delta\in[\Delta]$ in $G^{\prime}$. Since every node $v\in V$ has multiple representatives in $V^{\prime}$ it is necessary to ensure that a path found in $G^{\prime}$ does not contain cycles in $G$. Constraint (3f) ensures that at most one representative of any node $v\in V$ can be part of a path in $G^{\prime}$.
$\min\qquad$ $\sum_{j\in[N]}\sum_{\delta\in\Delta}\bm{c}_{j}^{\top}\bm{x}_{j}^{\delta}$ (3a)
s.t. $\bm{\ell}\in T(\bm{\phi}_{I}^{1},\ldots,\bm{\phi}_{I}^{N})$ (3b)
@@
$\sum_{f\in[F_{S}]}y^{\delta fk}\leq 1$ $\forall\delta\in[\Delta],k\in[K]$ (3g)
@@
$x_{j,e}^{\delta-1}\leq(1-\sum_{f\in[F_{S}]}y^{(\delta+1)fk})+(1-\ell_{j}^{k})$ $\forall\delta\in[\Delta-1],k\in[K],j\in[N],e\in E_{-\rightarrow}^{t\delta}$ (3i)
$x_{j,e}^{\delta}\leq(1-\sum_{f\in[F_{S}]}y^{(\delta+1)fk})+(1-\ell_{j}^{k})$ $\forall\delta\in[\Delta-1],k\in[K],j\in[N],e\in E_{-\downarrow}^{t\delta}$ (3j)
@@
$y^{\delta fk}\in\{0,1\}$ $\forall\delta\in[\Delta],f\in[F_{S}],k\in[K].$ (3l)
It is also essential to couple the district/layer assignment to the edges traversing between the layers. To control these edges, we use the binary variables $y^{\delta fk}$. Edges within one layer are not bounded by them. When set to one, $y^{\delta fk}$ indicates that district $f$ is assigned to layer $\delta$ in leaf $k$. Constraints (3g) ensure that at most one feature is assigned to every layer $\delta$ in every leaf $k$. The Constraints (3h) restrict the usage of edges between layers for every scenario that is assigned to leaf $k$. An edge $(u_{\delta},w_{\delta+1})$ connecting two consecutive layers can only be used if $y^{\delta fk}=1:f={f}(u)$ and $y^{(\delta+1)fk}=1:f={f}(w)$. Those edges can only be used if the districts assigned to the start and end layers match the features of the start and end nodes of that edge. Constraints (3i) and (3j) ensure that the sink nodes $t_{\delta}$ can only be visited in layer $\Delta$ or a layer $\delta\in[\Delta-1]$ if the subsequent layer $\delta+1$ has no district assigned, indicating the end of the solution sequence.
Note that if there are leaves that have no assigned scenarios, the formulation does not enforce it to contain a feasible solution. In practice, if such a case happens, the tree can either be pruned in post-processing, or an arbitrary feasible meta-solution can be placed in this leaf.